# Excitation of Inertial Modes in 3D Simulations of Rotating Convection

### v1

@@ Abstract
Thermal convection in rotating stars and planets drives anisotropic turbulence and strong differential rotation, both capable of feeding energy into global oscillations. Using 3D simulations of rotating convection in spherical shells, we show that inertial modes–oscillations restored by the Coriolis force–emerge naturally in rotationally constrained turbulence, without imposing any external forcing. By varying the rotation rate at fixed thermal forcing, we find that coherent modes appear only when the convective Rossby number, the ratio of the rotation period to the convective turnover time, falls below about one-half, where rotation dominates the dynamics. These modes are mostly retrograde in the rotating frame, equatorially symmetric, and confined to mid and high latitudes, with discrete frequencies well below twice the rotation rate. At lower viscosities, or smaller Prandtl number, mode excitation becomes more efficient and a broader spectrum of inertial modes emerges. While the precise excitation mechanism remains uncertain, our results suggest that the modes are driven by shear instabilities rather than stochastic forcing by convection. We conclude that similar inertial modes are likely to exist in the interiors of giant planets and stars, though their low frequencies will make them difficult to detect.
@@ 1 Introduction
Turbulence driven by thermal convection is a ubiquitous property of the atmospheres and interiors of planets and stars. The influence of the Coriolis force further organizes the flow and, through angular momentum transport, establishes differential rotation, which in turn feeds back on the convective motions. This shear shapes the global circulation patterns, influencing magnetic field generation and the transport of heat and composition. Moreover, such large-scale shear flows could provide a source of free energy capable of exciting a variety of oscillation modes, either through direct instabilities or resonant coupling with existing motions (Greenspan1968).
@@
The most common oscillation modes in rotating flows are inertial modes–oscillations restored by the Coriolis force. They have been observed across a wide range of astrophysical and geophysical systems, from the solar convection zone (Loptien2018; Gizon2021) and the convective cores of intermediate-mass stars (Ouazzani2020; Saio2021), to Earth’s outer core (e.g., Aldridge1987; Aldridge1988) and the oceans (e.g., Fu1981; Niu2023; Khimchenko2024). In these contexts, they have been shown to be valuable diagnostic tools for core rotation, differential rotation, as well as key contributors to ocean mixing, and atmosphere-ocean coupling. In addition, inertial modes play a central role in shaping the orbital and rotational evolution of star-planet and planet-moon systems (e.g., Ogilvie2004; Wu2005; Barker2022).
@@
The properties of inertial modes are not fully-understood (see, e.g., LeBars2015). This is largely due to their strong sensitivity to the geometry and boundary conditions of the flow, and their coupling to nonlinear and dissipative processes, which are challenging to model and often unresolved in simulations. Yet, studies of rotating fluids in spherical geometry, ranging from early asymptotic analyses of the linearized fluid equations (e.g., Bryan1889; Rieutord1997; Zhang2001; Rieutord2001) to modern laboratory and numerical experiments of spherical Couette flow (fluid confined in a spherical shell whose inner and outer boundaries rotate rigidly with different angular velocities, see, e.g., Triana2011; Matsui2011; Barik2018; Barik2024), have revealed several robust features. To name a few, they are waves drifting in the azimuthal direction with frequencies restricted to $|\omega|\leq 2\Omega$, where $\Omega$ is the angular frequency of the body or container. Also, within confined geometries, these waves focus along narrow shear layers or attractors, leading to enhanced shear and dissipation.
@@
Regarding their excitation, it has been primarily attributed to wave–mean flow interactions. One proposed mechanism is over-reflection (kelley2010), in which incident inertial waves are amplified upon reflection from a shear layer, thereby extracting energy from the background differential rotation. Another is excitation at a “critical layer”, often associated with the viscous shear layer that develops along the tangent cylinder in a rotating spherical shell (Rieutord2012). In this case, if a region of the fluid co-rotates with the drift frequency of an inertial wave, resonant coupling can occur, leading to the amplification of the mode. Follow up studies in the spherical Couette scenario did not find evidence of such a critical layer onset (Hoff2016; Barik2018) but did find some evidence of an onset through a shear instability of the axisymmetric background flow (Barik2018).
@@
As described above, most studies of inertial modes have focused on spherical Couette setups, where the imposed differential rotation is simple and well controlled by the boundary conditions of the inner and outer shells. In more realistic astrophysical and geophysical systems, differential rotation emerges self-consistently from rotating convection. In this context, only a few works (Bekki2022; Blume2024)–primarily focused on Rossby waves in the Sun–have identified inertial modes in nonlinear simulations of solar-like convection. These studies show that different families of inertial modes exist in both the radiative and convective zones of the Sun, each with distinct spatial structures and properties that highlight the complex interplay between differential rotation and convection in exciting inertial modes.
@@
In this work, we present a suite of 3D simulations of thermal convection in rotating spherical shells and analyze the resulting spatial and frequency spectra of the turbulence, along with the inertial modes excited in the system. We focus on how changes in the shell’s rotation rate affect mode properties when the strength of convective thermal driving is kept constant. Section 2 describes the model setup and numerical methods employed. Section 3 presents our analysis and main findings. Finally, in Section 4, we summarize our results and discuss them in light of previous studies, as well as their implications for real astrophysical objects.
2 Numerical Simulations
2.1 Hydrodynamical Model
@@
We simulate thermal convection in a 3D spherical shell of inner radius $r_{i}$ and outer radius $r_{o}$, rotating with constant angular frequency $\Omega_{0}{\mbox{{$\hat{z}$}}}$. The spherical shell boundaries are assumed to be isothermal, impenetrable, and stress-free. The kinematic viscosity $\nu$ and the thermal diffusivity $\kappa_{T}$ are assumed to be constant. For simplicity, we adopt the Boussinesq approximation (Spiegel_Veronis_1960), under which the flow is approximately incompressible, and density fluctuations are assumed to be small and linearly dependent on the temperature fluctuations, $\rho/\rho_{0}=-\alpha T$, where $\rho_{0}$ is the mean density of the fluid layer, and $\alpha$ is the coefficient of thermal expansion.
@@
We present the Boussinesq fluid equations in nondimensional form, using the shell depth $D=r_{o}-r_{i}$ as the unit of length and the viscous diffusion time $\tau_{\nu}=D^{2}/\nu$ as the unit of time. For the units of temperature, we use the temperature contrasts across the shell $\Delta T$. Finally, we adopt $\rho_{0}D^{2}/\tau_{\nu}^{2}$ as the unit of pressure. After nondimensionalization, the radial domain goes from $r_{i}=3$ to $r_{o}=4$, the shell depth is 1, and the resulting fluid equations become
@@
Figure 1: 3D snapshots of the radial velocity $u_{r}$. The velocity in simulations of rotating flows is normalized by $\Omega_{0}r_{o}$, while the non-rotating model is normalized by the free fall velocity $u_{\rm ff}$. Red and blue denote upflows and downflows, respectively. In the rotating cases, the flow exhibits markedly smaller, anisotropic spatial scales compared with the non-rotating case, where convection is isotropic. All the simulations have the same Rayleigh number $\mathrm{Ra}=5\times 10^{6}$ and Prandtl number $\mathrm{Pr}=1$, while the Ekman number varies from $\mathrm{Ek}\sim 7\times 10^{-5}$ to $\mathrm{Ek}=\infty$ (non-rotating case).
@@
where $\tau_{\kappa}=D^{2}/\kappa_{T}$ is the thermal diffusion time across the shell, $\tau_{\Omega}=1/2\Omega_{0}$ is the rotational timescale, and $\tau_{\rm ff}=\sqrt{D/\alpha g_{o}\Delta T}$ is the convective free-fall time across the shell, where $g_{o}$ is the acceleration of gravity at $r_{o}$.
@@
Systems dominated by rotation have low values of $\mathrm{Ro_{c}}$, while flows that are relatively insensitive to rotation have high $\mathrm{Ro_{c}}$. Our suite of simulations were conducted with the same Rayleigh and Prandtl numbers ($\mathrm{Ra}=5\times 10^{6}$, $\mathrm{Pr}=1$), but at different Ekman numbers ($\mathrm{Ek}\sim 7\times 10^{-5}\infty$). In terms of the convective Rossby number, our simulations have a broad dynamical range, with $\mathrm{Ro_{c}}\sim 0.17\infty$, thereby allowing us to explore both rotationally constrained and rotationally unconstrained convective regimes. For comparison, we also conduct a single simulation using $\mathrm{Pr}=0.1$, fixing $\mathrm{Ra}=5\times 10^{6}$ and $\mathrm{Ro_{c}}\approx 0.17$, so that $\mathrm{Ek}\sim 2.5\times 10^{-5}$. We present and analyze this simulation in detail in Section 3.5. We emphasize that realistic astrophysical parameters remain far beyond the reach of current computational capabilities. Nonetheless, by fixing $\mathrm{Ra}\gg 1,~\mathrm{Ek}\ll 1$, and $\mathrm{Pr}\lesssim 1$, we ensure that our simulations remain qualitatively within the same dynamical regime as those relevant to gas and ice giant planets and stars. For more details on the input parameters and the actual flow parameters achieved in the simulations, see Table 1.
Table 1: Simulation Parameters
Model Inputs Outputs
Ek $\mathrm{Ro_{c}}$ $\mathrm{Ro}_{r}$ $\mathrm{Ro}_{\theta}$ $\mathrm{Ro}_{\phi}$ $\mathrm{Re}_{r}$ $\mathrm{Re}_{\theta}$ $\mathrm{Re}_{\phi}$
$\infty$ $\infty$ $\infty$ $\infty$ $\infty$ 324 378 392
6.33 $\times$ 10-4 1.40 0.185 0.200 0.35 292 317 566
3.16 $\times$ 10-4 0.70 0.077 0.081 0.48 244 257 1534
2.37 $\times$ 10-4 0.53 0.048 0.053 0.40 203 223 1698
1.58 $\times$ 10-4 0.35 0.021 0.034 0.11 134 214 720
1.18 $\times$ 10-4 0.26 0.012 0.018 0.07 100 150 629
0.79 $\times$ 10-4 0.17 0.005 0.007 0.03 61 91 440
2.50 $\times$ 10-5 0.17 0.006 0.012 0.06 251 475 2292
Table 1: Input and output parameters for each model in this study. With the exception of the run in the last row, which has $\mathrm{Pr}=0.1$, all the rest of the simulations have the same Rayleigh number $\mathrm{Ra}=5\times 10^{6}$ and Prandtl number $\mathrm{Pr}=1$. The output Rossby numbers for each component of the velocity are calculated as $\mathrm{Ro}_{x}=u_{x,\rm rms}\mathrm{Ek}$, where $x$ denotes the radial, latitudinal, or longitudinal velocity. With our normalization, the output Reynolds numbers for each component are simply $\mathrm{Re}_{x}=u_{x,\rm rms}$.
2.2 Numerical Methods
@@
3 Results
3.1 Morphology and Spatial Spectra of the Flow
@@
The radial velocity (Figure 1) illustrates the morphological changes of the convective flow as rotation is varied. In the non-rotating case, the flow is isotropic, with broad upwellings and downwellings and no preferred horizontal scale or alignment. As rotation increases (decreasing $\mathrm{Ro_{c}}$), the convective structures narrow and align with the rotation axis, forming the columnar patterns characteristic of rapidly rotating convection. At the smallest $\mathrm{Ro_{c}}$, the anisotropy between vertical and horizontal scales is most pronounced, and the flow is strongly constrained by the Coriolis force.
The corresponding azimuthal velocity (Figure 2) reveal the large-scale zonal flows that develop in rotating convection. These flows arise from Reynolds stresses generated by the tilt of convective columns under rotation. For the most rapidly rotating cases ($\mathrm{Ro_{c}}\lesssim 0.5)$, a strong prograde (eastward) jet forms at the equator, flanked by retrograde (westward) flows at higher latitudes. As $\mathrm{Ro_{c}}$ increases, the amplitude and structure of these jets change: the equatorial prograde flow weakens, reverses sign, and gives way to a retrograde jet, with prograde flows at higher latitudes. This transition is consistent with previous studies linking the direction of equatorial jets to the convective Rossby number (the well known solar to anti-solar differential rotation, see, e.g., Gastine2014; Camisassa2022). In the absence of rotation, the zonal component lacks any coherent structure, and differential rotation does not develop.
The influence of rotation on convection is further reflected in differences in the corresponding spatial power spectra. We compute these spectra using the SHTns package (Schaeffer2013), which employs a vector spherical harmonic decomposition of the velocity field at a given radius $r$
@@
$\mathcal{P}(\ell)=\sum_{m\geq 0}C_{m}\left(|Q_{\ell}^{m}|^{2}+\ell(\ell+1)(|S_{\ell}^{m}|^{2}+|T_{\ell}^{m}|^{2})\right)~,$ (8)
@@
Figure 3: Velocity power spectra for all simulations. Flows were sampled at a radius $r=0.85r_{o}$ to construct the spectra, with other radii yielding essentially identical results. Panels (a) and (b) show the power associated with all azimuthal wavenumbers $m$ and with only the non-axisymmetric components ($m>0$), respectively. Panel (c) shows the power of the zonal differential rotation, i.e., the power considering only the toroidal component with the $m=0$ contribution in Equation (8). The dashed lines in panels (b) and (c) have a slope of 2/3-2/3 and 5-5 for comparison with Kolmogorov spectrum (Kolmogorov1941) and with zonostropic turbulence (Boning2023), respectively.
Figure 3 shows the power spectra for all our simulations, computed using the full set of azimuthal wavenumbers $m$ and, separately, using only the non-axisymmetric components ($m>0$), which represent the convective turbulence. All spectra were calculated at radius $r=0.85r_{o}$ and temporally averaged over a viscous diffusion time. We emphasize that the results are not sensitive to the choice of radius, since the flow is nearly incompressible and density variations are small across the shell depth. In the absence of rotation, power simply decreases from large to small scales, as expected for non-rotating convection. With rotation, the distribution is more structured, axisymmetric motions ($m=0$) dominate at low degrees and show peaks at odd $\ell$ values arising from toroidal equatorially symmetric zonal flows, while the convective part ($m>0$) grows with $\ell$ up to a critical value before declining in a way reminiscent of the non-rotating case.
@@
As argued by Rhines1975, the characteristic jet width $d$ and speed $U$ in planetary atmospheres satisfy $d\sim\sqrt{UR/2\Omega\sin\theta}$, where $\Omega$ is the rotation rate and $\theta$ the colatitude. Identifying the jet width with spherical harmonic degree, $d\sim\pi R/\ell$, and estimating the jet velocity from the definition of the zonal energy spectrum, $U\sim\sqrt{2\ell P_{\ell,\rm zonal}}$, which is dominated by low $\ell$ modes, one obtains the scaling $P_{\ell,\rm zonal}\propto\ell^{-5}$.
Figure 4: Power spectra of the kinetic energy at $r=0.85r_{o}$, as a function of frequency for all rotating models. We show the frequency normalized to $2\Omega_{0}$, and to the convective frequency $\omega_{\rm conv}$. The power is normalized by the total over all frequencies, $\ell$, and $m$. Black curves show the sum over all $m$ and $\ell$, while colored curves indicate the contributions from all $\ell$ but individual $m$, allowing identification of the modes responsible for the observed peaks. The vertical dashed line in each panel marks the frequency associated with the evolution of mean flows (“zero-frequency” differential rotation), set by the Ekman time of the rotating flow.
@@
3.2 Frequency Spectra
A direct way to identify the excitation of oscillation modes in a simulation is to analyze the flow in frequency space. We proceed in much the same way as when examining the flow across spatial scales, by constructing the power spectrum as a function of spherical harmonic degree $\ell$, order $m$, and frequency $\omega$. The starting point is recording of the velocity field at a fixed radius over a sequence of equally spaced time intervals. At each snapshot, every component of the velocity is expanded in spherical harmonics, yielding a time series of complex spectral coefficients $u^{r,\theta,\phi}_{\ell m}(t)$ for each ($\ell,m$) mode. Applying a temporal Fourier transform to each coefficient reveals its oscillatory content, and the squared amplitudes $|\hat{u}^{r,\theta,\phi}_{\ell m}(\omega)|^{2}$ measure the kinetic energy associated with each mode at a given frequency. Summing over all the spherical harmonic degree $\ell$, order $m$, and all velocity components, produces the total kinetic energy spectrum per frequency bin, $\mathcal{P}(\omega)\approx(dE/d\omega)\Delta\omega$

### v2

@@ Abstract
Thermal convection in rotating stars and planets drives anisotropic turbulence and differential rotation, both capable of feeding energy into global oscillations. Using 3D simulations of rotating convection in spherical shells, we show that inertial modes–oscillations restored by the Coriolis force–emerge naturally in rotationally constrained turbulence, without imposing any external forcing other than thermal/buoyancy driving. By varying the rotation rate at fixed Rayleigh number, we find that coherent modes appear only when the convective Rossby number, the ratio of the rotation period to the convective turnover time, falls below about one-half, where rotation dominates the dynamics. These modes are mostly retrograde in the rotating frame, equatorially symmetric, and confined to mid and high latitudes, with discrete frequencies well below twice the background rotation rate. At lower viscosities, or smaller Prandtl number, mode excitation becomes more efficient and a broader spectrum of inertial modes emerges. While the precise excitation mechanism remains uncertain, our results suggest that the modes are driven by instabilities due to differential rotation rather than stochastic forcing by convection. We conclude that similar inertial modes are likely to exist in the interiors of giant planets and stars, though their low frequencies will make them difficult to detect.
@@ I Introduction
Turbulence driven by thermal convection is a ubiquitous property of the atmospheres and interiors of planets and stars. The influence of the Coriolis force further organizes the flow and, through angular momentum transport, establishes differential rotation, which in turn feeds back on the convective motions. This shear shapes the global circulation patterns, influencing magnetic field generation and the transport of heat and composition. Moreover, such large-scale shear flows could provide a source of free energy capable of exciting a variety of oscillation modes, either through direct instabilities or through nonlinear self-interactions of inertial waves, including triadic resonances (Greenspan1968; Barik2018; Lin2021).
@@
The most common oscillation modes in rotating flows are inertial modes–oscillations restored by the Coriolis force. They have been observed across a wide range of astrophysical and geophysical systems, from the solar convection zone (Loptien2018; Gizon2021) and the convective cores of intermediate-mass stars (Ouazzani2020; Saio2021), to Earth’s outer core (e.g., Aldridge1987; Aldridge1988) and the oceans (e.g., Fu1981; Niu2023; Khimchenko2024). In these contexts, they have been shown to be valuable diagnostic tools for core rotation, differential rotation, as well as key contributors to ocean mixing, and atmosphere-ocean coupling. In addition, inertial modes play a central role in shaping the orbital and rotational evolution of stellar binaries, including star-planet and planet-moon systems (e.g., Ogilvie2004; Wu2005; Barker2022).
@@
The properties of inertial modes are not fully-understood (see, e.g., LeBars2015). This is largely due to their strong sensitivity to the geometry and boundary conditions of the flow, and their coupling to nonlinear and dissipative processes, which are challenging to model and often unresolved in simulations. Yet, studies of rotating fluids in spherical geometry, ranging from early analyses of the linearized fluid equations (e.g., Bryan1889; Rieutord1997; Zhang2001; Rieutord2001) to modern laboratory and numerical experiments of spherical Couette flow (fluid confined in a spherical shell whose inner and outer boundaries rotate rigidly with different angular velocities, see, e.g., Triana2011; Matsui2011; Barik2018; Barik2024), have revealed several robust features. To name a few, they are global waves with frequencies restricted to $|\omega|\leq 2\Omega$, where $\Omega$ is the angular frequency of the body or container, and non-axisymmetric modes drift azimuthally at a rate $\omega/m$, with $m$ the azimuthal wavenumber. In confined geometries, these waves focus along narrow shear layers or attractors, leading to enhanced shear and dissipation.
@@
Regarding their excitation, it has been primarily attributed to wave–mean flow interactions. One proposed mechanism is over-reflection (kelley2010), in which incident inertial waves are amplified upon reflection from a shear layer, thereby extracting energy from the background differential rotation. In rotating bodies deformed by tidal interactions, inertial modes can be excited through the elliptical instability, a parametric instability in which pairs of inertial waves interact with the tidally induced strain field and extract energy from the tidal flow (see, e.g., deVries2023). Another is excitation at a "critical layer", often associated with the viscous shear layer that develops along the tangent cylinder in a rotating spherical shell (Rieutord2012; Astoul2021). In this case, if a region of the fluid co-rotates with the drift frequency of an inertial wave, resonant coupling can occur, leading to the amplification of the mode (i.e., critical layers correspond to regions of corotation resonance). Follow up studies of spherical Couette flow found no evidence for excitation at critical layers (Hoff2016; Barik2018), but did find excitation via shear instabilities of the axisymmetric background flow (Barik2018).
@@
As described above, most studies of inertial modes have focused on spherical Couette setups, where the imposed differential rotation is simple and well controlled by the boundary conditions of the inner and outer shells. In more realistic astrophysical and geophysical systems, differential rotation emerges self-consistently from rotating convection. Inertial modes have been reported in only a handful of nonlinear simulations of rotating convection, notably in studies of Rossby waves in using spherical models of solar-like convection (Bekki2022; Blume2024) and in simulations examining mode excitation in both full spheres and Cartesian geometries (Lin2021; deVries2023). These studies have shown that different families of inertial modes exist in both the radiative and convective zones of stars and planets, each with distinct spatial structures and properties that highlight the complex interplay between differential rotation and convection in exciting inertial modes.
@@
In this work, we present a suite of 3D simulations of thermal convection in rotating spherical shells and analyze the resulting spatial and frequency spectra of the turbulence, along with the inertial modes excited in the system. We focus on how changes in the shell’s rotation rate affect mode properties when the strength of convective thermal driving is kept constant. Section II describes the model setup and numerical methods employed. Section III presents our analysis and main findings. Finally, in Section IV, we summarize our results and discuss them in light of previous studies, as well as their implications for real astrophysical objects.
II Numerical Simulations
II.1 Hydrodynamical Model
@@
We simulate thermal convection in a 3D spherical shell of inner radius $r_{i}$ and outer radius $r_{o}$, initially rotating with constant angular frequency $\Omega_{0}{\mbox{{$\hat{z}$}}}$. The gravity profile is ${\mbox{{$g$}}}=g_{0}(r/r_{o})$. The spherical shell boundaries are assumed to be isothermal, impenetrable, and stress-free. The kinematic viscosity $\nu$ and the thermal diffusivity $\kappa_{T}$ are assumed to be constant. For simplicity, we adopt the Boussinesq approximation (Spiegel_Veronis_1960), under which the flow is approximately incompressible, and density fluctuations are assumed to be small and linearly dependent on the temperature fluctuations, $\rho/\rho_{0}=-\alpha T$, where $\rho_{0}$ is the mean density of the fluid layer, $\alpha$ is the coefficient of thermal expansion, and $T$ is the the temperature perturbation relative to a constant reference temperature $T_{0}$.
@@
We present the Boussinesq fluid equations in nondimensional form, using the shell depth $D=r_{o}-r_{i}$ as the unit of length and the viscous diffusion time $\tau_{\nu}=D^{2}/\nu$ as the unit of time. For the units of temperature, we use the temperature contrasts across the shell $\Delta T$. Finally, we adopt $\rho_{0}D^{2}/\tau_{\nu}^{2}$ as the unit of pressure. After nondimensionalization, the radial domain goes from $r_{i}=3$ to $r_{o}=4$, i.e., the fractional radius is $r_{i}/r_{o}=0.75$, the shell depth is 1, and the resulting fluid equations become
@@
Table 1: Input and output parameters for each model in this study. With the exception of the last row, which has $\mathrm{Pr}=0.1$, all simulations use $\mathrm{Ra}=5\times 10^{6}$ and $\mathrm{Pr}=1$. $\mathrm{Ra_{c}}$ denotes the critical Rayleigh number for the onset of convection at a given Ekman number. The first row corresponds to a non-rotating reference case. In our nondimensionalization, output Rossby numbers are defined as $\mathrm{Ro}_{x}=u_{x,\mathrm{rms}}\mathrm{Ek}$, where $x=r,\theta,\phi$, while the corresponding Reynolds numbers reduce to the dimensionless rms velocity, $\mathrm{Re}_{x}=u_{x,\mathrm{rms}}$.
Model inputs Outputs
Ek $\mathrm{Ro_{c}}$ $\mathrm{Ra_{c}}$ $\mathrm{Ro}_{r}$ $\mathrm{Ro}_{\theta}$ $\mathrm{Ro}_{\phi}$ $\mathrm{Re}_{r}$ $\mathrm{Re}_{\theta}$ $\mathrm{Re}_{\phi}$
$\infty$ $\infty$ $3.0\times 10^{3}$ $\infty$ $\infty$ $\infty$ 324 378 392
$6.33\times 10^{-4}$ 1.40 $2.0\times 10^{4}$ 0.185 0.200 0.35 292 317 566
$3.16\times 10^{-4}$ 0.70 $4.7\times 10^{4}$ 0.077 0.081 0.48 244 257 1534
$2.37\times 10^{-4}$ 0.53 $6.7\times 10^{4}$ 0.048 0.053 0.40 203 223 1698
$1.58\times 10^{-4}$ 0.35 $1.1\times 10^{5}$ 0.021 0.034 0.11 134 214 720
$1.18\times 10^{-4}$ 0.26 $1.6\times 10^{5}$ 0.012 0.018 0.07 100 150 629
$0.79\times 10^{-4}$ 0.17 $2.6\times 10^{5}$ 0.005 0.007 0.03 61 91 440
$2.50\times 10^{-5}$ 0.17 $3.0\times 10^{5}$ 0.006 0.012 0.06 251 475 2292
@@
where $\tau_{\kappa}=D^{2}/\kappa_{T}$ is the thermal diffusion time across the shell, $\tau_{\Omega}=1/2\Omega_{0}$ is the rotational timescale, and $\tau_{\rm ff}=\sqrt{D/\alpha g_{o}\Delta T}$ is the convective free-fall time across the shell, where $g_{o}$ is the acceleration of gravity at $r_{o}$. Another dimensionless number that characterize the flows is the Reynolds number, which in terms of the free fall velocity $u_{\rm ff}=\sqrt{\alpha g_{o}\Delta TD}$ can be written as a combination of the Rayleigh and Prandtl numbers, $\mathrm{Re}=u_{\rm ff}D/\nu=(\mathrm{Ra}/\mathrm{Pr})^{1/2}$.
@@
Systems dominated by rotation have low values of $\mathrm{Ro_{c}}$, while flows that are relatively insensitive to rotation have high $\mathrm{Ro_{c}}$. Our suite of simulations were conducted with the same Rayleigh and Prandtl numbers ($\mathrm{Ra}=5\times 10^{6}$, $\mathrm{Pr}=1$), but at different Ekman numbers ($\mathrm{Ek}\sim 7\times 10^{-5}\infty$). In terms of the convective Rossby number, our simulations have a broad dynamical range, with $\mathrm{Ro_{c}}\sim 0.17\infty$, thereby allowing us to explore both rotationally constrained and rotationally unconstrained convective regimes. For comparison, we also conduct a single simulation using $\mathrm{Pr}=0.1$, fixing $\mathrm{Ra}=5\times 10^{6}$ and $\mathrm{Ro_{c}}\approx 0.17$, so that $\mathrm{Ek}\sim 2.5\times 10^{-5}$. We present and analyze this simulation in detail in Section III.5.
All our rotating simulations are conducted at supercriticalities $\mathrm{Ra}/\mathrm{Ra_{c}(Ek)}$ in the rage 10–260, where $\mathrm{Ra_{c}(Ek)}$ is the critical Rayleigh number for the onset of convection at a given Ekman number. The values of the critical Rayleigh numbers $\mathrm{Ra_{c}}$ were obtained by interpolating values in the database of BarikEtAl2023 except for the $\mathrm{Pr}=0,1$ case, which was computed using the linear code Kore (https://github.com/repepo/kore, BarikEtAl2023). The non-rotating simulation has $\mathrm{Ra_{c}}\approx 3\times 10^{3}$, so that $\mathrm{Ra}/\mathrm{Ra_{c}}\approx 1660$. We also emphasize that realistic astrophysical parameters remain far beyond the reach of current computational capabilities. Nonetheless, by fixing $\mathrm{Ra}\gg 1,~\mathrm{Ek}\ll 1$, and $\mathrm{Pr}\lesssim 1$, we ensure that our simulations remain qualitatively within the same dynamical regime as those relevant to gas and ice giant planets and stars. For more details on the input parameters and the actual flow parameters achieved in the simulations, see Table 1.
II.2 Numerical Methods
@@
III Results
III.1 Morphology and Spatial Spectra of the Flow
Figure 1: 3D snapshots of the radial velocity $u_{r}$. The velocity in simulations of rotating flows is normalized by $\Omega_{0}r_{o}$, while the non-rotating model is normalized by the free fall velocity $u_{\rm ff}$. Red and blue denote upflows and downflows, respectively. In the rotating cases, the flow exhibits markedly smaller, anisotropic spatial scales compared with the non-rotating case, where convection is more isotropic. All the simulations have the same Rayleigh number $\mathrm{Ra}=5\times 10^{6}$ and Prandtl number $\mathrm{Pr}=1$, while the Ekman number varies from $\mathrm{Ek}\sim 7\times 10^{-5}$ to $\mathrm{Ek}=\infty$ (non-rotating case).
@@ III.1 Morphology and Spatial Spectra of the Flow
The radial velocity (Figure 1) illustrates the morphological changes of the convective flow as rotation is varied. In the non-rotating case, the flow is more isotropic, with broad upwellings and downwellings and no preferred horizontal scale or alignment. As rotation increases (decreasing $\mathrm{Ro_{c}}$), the convective structures narrow and align with the rotation axis, forming the columnar patterns characteristic of rapidly rotating convection. At the smallest $\mathrm{Ro_{c}}$, the anisotropy between vertical and horizontal scales is most pronounced, and the flow is strongly constrained by the Coriolis force.
The corresponding azimuthal velocity (Figure 2) reveal the large-scale zonal flows that develop in rotating convection. These flows arise from Reynolds stresses generated in the convection zone (see, e.g., Christensen2001; Busse2002; Aurnou2007). For the most rapidly rotating cases ($\mathrm{Ro_{c}}\lesssim 0.5)$, a strong prograde (eastward) jet forms at the equator, flanked by retrograde (westward) flows at higher latitudes. As $\mathrm{Ro_{c}}$ increases, the amplitude and structure of these jets change: the equatorial prograde flow weakens, reverses sign, and gives way to a retrograde jet, with prograde flows at higher latitudes. This transition is consistent with previous studies linking the direction of equatorial jets to the convective Rossby number (the well known solar to anti-solar differential rotation, see, e.g., Gastine2014; Camisassa2022). In the absence of rotation, the zonal component lacks any coherent large-scale structure, and no mean (axisymmetric) differential rotation develops.
The influence of rotation on convection is further reflected in differences in the corresponding spatial power spectra. We compute these spectra using the SHTns package (Schaeffer2013), which employs a vector spherical harmonic decomposition of the velocity field at a given radius $r$ and time $t$
@@
$\mathcal{P}(\ell,r,t)=\sum_{m\geq 0}C_{m}\left(|Q_{\ell}^{m}|^{2}+\ell(\ell+1)(|S_{\ell}^{m}|^{2}+|T_{\ell}^{m}|^{2})\right)~,$ (8)
@@
Figure 3: Velocity power spectra for all simulations. Flows were sampled at a radius $r=0.85r_{o}$ to construct the spectra, with other radii yielding essentially identical results. Panels (a) and (b) show the power associated with all azimuthal wavenumbers $m$ and with only the non-axisymmetric components ($m>0$), respectively. Panel (c) shows the power of the zonal flow, i.e., the power considering only the toroidal component with the $m=0$ contribution in Equation (8). The dashed lines in panels (b) and (c) have a slope of 2/3-2/3 and 5-5 for comparison with Kolmogorov spectrum (Kolmogorov1941) and with zonostropic turbulence (Boning2023), respectively.
Figure 3 shows the power spectra for all our simulations, computed using the full set of azimuthal wavenumbers $m$ and, separately, using only the non-axisymmetric components ($m>0$), which represent the convective turbulence. All spectra were calculated at radius $r=0.85r_{o}$ and temporally averaged over a viscous diffusion time. We emphasize that the results are not sensitive to the choice of radius within the bulk of the shell, since the flow is nearly incompressible and density variations are small across the shell depth. In the absence of rotation, power simply decreases from large to small scales, as expected for non-rotating convection. With rotation, the distribution is more structured, axisymmetric motions ($m=0$) dominate at low degrees and show peaks at odd $\ell$ values arising from toroidal equatorially symmetric zonal flows, while the convective part ($m>0$) grows with $\ell$ up to a critical value before declining in a way reminiscent of the non-rotating case.
@@
As argued by Rhines1975 in the context of $\beta$-plane turbulence, the characteristic jet width $d$ and speed $U$ in planetary atmospheres satisfy $d\sim\sqrt{UR/2\Omega\sin\theta}$, where $\Omega$ is the rotation rate and $\theta$ the colatitude. Adopting this scaling and extrapolating it to spherical geometry, we identify the jet width with spherical harmonic degree, $d\sim\pi R/\ell$, and estimate the jet velocity from the definition of the zonal energy spectrum, $U\sim\sqrt{2\ell P_{\ell,\rm zonal}}$. Since the zonal spectrum is dominated by low-$\ell$ modes, one obtains the scaling $P_{\ell,\rm zonal}\propto\ell^{-5}$.
Figure 4: Power spectra of the kinetic energy at $r=0.85r_{o}$, as a function of frequency for all rotating models. We show the frequency normalized to $2\Omega_{0}$, and to the convective frequency $\omega_{\rm ff}$. The power is normalized by the total over all frequencies, $\ell$, and $m$. Black curves show the sum over all $m$ and $\ell$, while colored curves indicate the contributions from all $\ell$ but individual $m$, allowing identification of the modes responsible for the observed peaks.
@@
III.2 Frequency Spectra
A direct way to identify the excitation of oscillation modes in a simulation is to analyze the flow in frequency space. We proceed in much the same way as when examining the flow across spatial scales, by constructing the power spectrum as a function of spherical harmonic degree $\ell$, order $m$, and frequency $\omega$. The starting point is recording of the velocity field at a fixed radius over a sequence of equally spaced time intervals. At each snapshot, every component of the velocity is expanded in spherical harmonics, yielding a time series of complex spectral coefficients $u^{r,\theta,\phi}_{\ell m}(t)$ for each ($\ell,m$) mode. Applying a temporal Fourier transform to each coefficient reveals its oscillatory content, and the squared amplitudes $|\hat{u}^{r,\theta,\phi}_{\ell m}(\omega)|^{2}$ measure the kinetic energy associated with each mode at a given frequency. Summing over all the spherical harmonic degree $\ell$, order $m$, and all velocity components, produces the total kinetic energy spectrum per frequency bin, $\mathcal{P}(\omega)\approx(dE/d\omega)\Delta\omega$,